# Deviations from the Nash equilibrium and emergence of tacit collusion in a two-player optimal execution game with reinforcement learning

### v1

@@ 3.1.1 Action selection and reward function
For each of the $N$ time-steps and $C$ trading iterations, the agents’ knowledge of the state of market environment is the tuple $g^{i}_{t}=(t,q^{i}_{t},S_{t-1})$ for $i=1,2$. Thus, each agent knows the current time-step, her individual remaining inventory, and the permanently impacted mid-price at the previous time step. Clearly, only the first and the last are common knowledge of the two agents, while no information on inventory $q_{t}$ or past selling actions $v_{t}$ is shared between the agents.
@@
Then, depending on the current value of $\epsilon$, a draw $\zeta$, different for each agent, from a uniform distribution determines whether the agent performs exploration or exploitation. This means that with probability $\epsilon$ the agent chooses to explore and thus she chooses at time-step $t$ the quantity to sell $v_{t}$ sampling from a normal distribution with mean $\mu=\frac{q_{t}}{N-t}$ and standard deviation $\delta=|\frac{q_{t}}{N-t}|$. In this way, on average we favour a TWAP execution, allowing for both positive and negative values of $v_{t}$, meaning that sell and buy actions can be selected. In the exploration phase we bind the quantity to be sold to be $v_{i,t}\in[-q_{0},q_{0}],\,\,\,\,i=1,2$. Alternatively, with probability $1-\epsilon$, the agent chooses the optimal Q-action as the one that maximises the Q-values from $Q_{\text{main}}$ net, thus exploiting what learnt in the exploration phase. We bind the agents to sell all their inventory within the considered time window, still exploring a large number of states, rewards, and actions.
@@
Once every $m$ actions taken during training iterations by both the agents, $\epsilon$ is multiplied by a constant $c<1$ such that $\epsilon\to 0$ as $m\to\infty$. In this way, for a large number of iterations $C$, $\epsilon$ converges to zero and the algorithm gradually stops exploring and starts to greedily exploit what the agent has learned in terms of weights $\theta$ of the $Q_{\text{main}}$ net. Notice that the update for $\epsilon$ happens at the same rate for both the agents, thus they explore and exploit contemporaneously, while the draw from the uniform distribution $\zeta$ is different for every agent. For each time step $t$ we decide which of the two agents trades first with a coin toss. Once that the ordering of the trades is decided, the action decision rule in the training phase unfolds as:
@@
$\begin{split}&\epsilon\in(0,1)\,\,,\,\,\zeta\sim\mathcal{U}(0,1)\\ &v_{t}=\begin{cases}\sim\mathcal{N}(\mu=\frac{q_{t}}{N-t},\delta=|\frac{q_{t}}{N-t}|)\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,,\text{if}\,\,\zeta\leq\epsilon\\ \operatorname*{arg\,max}_{v^{\prime}\in[0,q_{t}]}Q_{M}(g^{i}_{t},v^{\prime}|\theta_{\text{main}})\,\,\,\,\,\,\,\,,\text{else}\end{cases}\end{split}$ (22)
@@
$\begin{split}r_{1,i}&=S_{0}v_{1,i}-\alpha v_{1,i}^{2}\\ r_{2,i}&=(S_{1}+\kappa(v_{1,i}+v_{1,i^{-}})+\xi)v_{2,i}-\alpha v_{2,i}^{2}\\ \vdots&\\ r_{N,i}&=(S_{N-1}+\kappa(v_{N-1,i}+v_{N-1,i^{-}})+\xi)v_{N,i}-\alpha v_{N,i}^{2}\\ \end{split}$ (24)
@@
Thus, for each time step $t$ the agent $i$ sees the reward from selling $v_{t,i}$ shares, and stores the state of the environment, $g_{t,i}$, where the sell action was chosen along with the reward and the next state $g_{t+1,i}$ where the environment evolves. At the end of the episode, the reward per episode per agent is $-q_{0}S_{0}+\sum^{N}_{t=1}S_{t-1}v_{t,i}-\alpha v_{t,i}^{2}$. Written in this form, the aim of the agent is to cumulatively maximise such reward, such that the liquidation value of the inventory happens as close as possible to the initial value of the portfolio.
@@ 3.1.2 Training scheme
NN layers $5$ $C$ train its. $5,000$ $N$ intervals $10$
Hidden nodes $30$ $M$ test its. $2,500$ $S_{0}$ price $10\$$
ADAM lr $0.0001$ $m$ reset rate $75$ acts. $q_{0}$ inventory $100$
@@
$L$ mem. len. $15,000$ $\gamma$ discount $1$ $\kappa$ p. impact $0.001$
Bearing in mind that the procedure is exactly the same for both agents, we focus now on the training scheme of one of them, dropping the subscript $i$. We let the states $g_{t}$, actions $v_{t}$, rewards $r_{t}$ and subsequent future states $g_{t+1}$ obtained by selling a quantity $v_{t}$ in state $g_{t}$, to form a transition tuple that is stored into a memory of maximum length $L$, we have two different memories, one for each agent. As soon as the memory contains at least $b$ transitions, the algorithm starts to train the Q-nets. To this end, the algorithm samples random batches of length $b$ from the memory of the individual agent, and for each sampled transition $j$ it individually calculates:
$y^{j}_{t}(\theta_{\text{tgt}})=\begin{cases}r^{j}_{t}&\text{if }{t}=N;\\ r^{j}_{t}+\gamma Q_{\text{tgt}}(g^{j}_{t+1},v^{*}|\theta_{\text{tgt}})&\text{else}\end{cases}$ (25)
In Eq. (25), $r^{j}_{t}$ is the reward for the time step considered in the transition $j$, $g^{j}_{t+1}$ is the subsequent state reached in $t+1$, known since it is stored in the same transition $j$, while $v^{*}=\operatorname*{arg\,max}_{v}Q_{\text{main}}(g^{j}_{t},v|\theta_{\text{main}})$. $\gamma$ is a discount factor that accounts for risk-aversion [^5] We set a $\gamma=1$ since we model risk neutral agents. Each agent individually minimises the mean squared error loss between the target $y(\theta_{\text{tgt}})$ and the values obtained via the $Q_{\text{main}}$ net. In formulae:
@@
The features of the Q-nets for each agent are ($q_{i,t},t,S_{t-1},v_{i,t}$) and are normalised in the domain $[-1,1]$ using the procedure suggested in [ning2021double], whereas normalised mid-prices $\bar{S}_{t}$ are obtained via min-max normalisation. In our setup we use fully connected feed-forward neural networks with $5$ layers, each with $30$ hidden nodes. The activation functions are leakyReLu, and finally we use ADAM for optimisation. With the exception of the volatility parameter $\sigma$ which will be specified later, the parameters used in the algorithm are reported in Table 1, the training algorithm is reported in Algorithm 1.
@@
                  $v_{i,t}\leftarrow\begin{cases}\text{sample}\,\,\,\mathcal{N}(\mu=\frac{q_{t}}{N-t},\delta=|\frac{q_{t}}{N-t}|)\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,,\text{with probability $\epsilon$}\\ \operatorname*{arg\,max}_{v^{\prime}\in[0,q_{t}]}Q_{M}(g_{t},v^{\prime}|\theta_{\text{main}})\,\,\,\,\,\,\,\,,\text{with probability ($1-\epsilon$)}\end{cases};$
@@
Figure 1: Scatter plot of the IS of the two agents for 20 testing runs of 2500 iterations in the zero noise case ($\sigma=10^{-9}$).
@@ 4 Results
In this Section we present the results obtained by using the RL Algorithm 1 in the game setting outlined in Section 2. The experiments aim at studying the existence and the form of the learnt equilibria, by analysing the policies chosen by the two interacting trading agents. We then compare the equilibria with the Nash equilibrium, the Pareto-efficient set of solutions, and the Pareto optimal strategy. We consider different scenarios, using the market structure as outlined in Eq. (4).
It can be easily noticed how the theoretical Nash equilibrium for this game does not depend on the volatility level of the asset. In fact, when the risk aversion parameter $\lambda=0$, the Nash equilibrium of Eq. (7) depends only on the permanent and temporary impact coefficients $\kappa$ and $\alpha$, respectively. However, in the numerical determination of the equilibrium solution, volatility $\sigma$ and the associated diffusion play the role of a disturbance term, since an agent cannot determine whether an observed price change is only due to the impact generated by the trading of the other agent or if it is a random fluctuation due to volatility. In this sense, volatility plays the role of a noise-to-signal ratio here. To quantify the effect of volatility on the learnt equilibria, we perform three sets of experiments for different levels of the volatility parameter $\sigma$, leaving unchanged the impact parameters. More specifically we consider, for both training and testing phase the case where $\sigma=10^{-9}$ (termed zero noise case), $\sigma=10^{-3}$ (moderate noise) and $\sigma=10^{-2}$ (large noise) case. In all three cases, we use the same temporary and permanent impacts $\alpha=0.002$ and $\kappa=0.001$.
@@ 4.1 The zero noise case
The results of the experiments are displayed in Figure 1, which shows, as a scatter plot, the $IS$ of both agents in each iteration of the testing phase. Each colour represents the results of one of the $20$ runs of $2,500$ iterations each. The blue circles shows the $IS$ centroids per testing run (the number in the circle identifies the run). For comparison the plot reports as a red star the point corresponding to the Nash equilibrium of Eq. (7), which is used to divide the graph into four quadrants, delimited by the red dashed lines. The top-right quadrant contains point which are sub-optimal with respect to the Nash equilibrium for both players, while the points in the top-left and bottom-right quadrant report points where the found equilibrium favours one of the two agents at the expenses of the other. In particular, in these regions one of the agents achieve an IS smaller than the one of the Nash equilibrium, while the other agents performs worst. In a sense, one of the agents predates on the other in terms of reward. The bottom-left quadrant is the most interesting, since here both agents are able to achieve an $IS$ smaller than the one in the Nash equilibrium and thus contain potential collusive equilibria. For comparison, the black star indicates the $IS$ of the Pareto-optimal strategy found in Theorem 2, while the magenta line is the linear combination of the Pareto-optimal $IS$. Finally, the green rectangle denotes the area between the Nash and the Pareto-optimal $IS$s, and we call it the collusive area. In fact, in that area we find the $IS$s for both the agents that lie between the proper collusion defined by the Pareto-optimum and the Nash equilibrium.
@@
Figure 2: Optimal execution strategies for 20 testing runs of 2500 iterations, using $\sigma=10^{-9}$.
@@
Looking at Figure 1 we can notice that the $IS$s per testing run concentrate in the collusive area of the graph (green rectangle) between the Nash and the Pareto-optimal $IS$. This means that when the noise is minimal, it is easier for the agents to adopt tacit collusive behaviour and obtain costs that fall near the collusion $IS$, i.e. the Pareto-optimum. The majority of the remaining $IS$s per iteration still lie within the second and fourth quadrant of the graph, meaning that again they are in general able to find strategies that, at each iteration, do not allow one agent to be better without worsening the other.
@@
Considering the strategies found by the agents in Figure 2, we notice how the agents keep on trading at different speed, i.e. their selling policies are consistent with the presence of a slow trader and a fast trader. It can be seen how the policies followed by the agents in almost all the 20 simulations depend inversely on the strategy adopted by the competitor. Moreover, in most of the runs in the collusive region, the average strategy of the agents is very similar to the TWAP strategy. This is possible thanks to very low noise, and in this case the agents are able to find more easily the Pareto-optimal strategy, which corresponds to an equilibrium where they both pay the lowest amount possible in terms of $IS$ by tacitly colluding with their trading.
@@
These evidences underline the quite intuitive fact that, when the noise to signal ratio is low it is simple for the agents to disentangle their actions from those of other agents. The agents find an incentive to deviate from the Nash-equilibrium adopting strategies that allow for lower costs. The majority of these strategies correspond to an average collusive $IS$s, thus with lower cost than in the Nash equilibrium, but still slightly greater than the Pareto-optimum in terms of $IS$s. In general, per iteration neither agent can be better off without increasing the other agent cost level. There is thus a strong evidence of tacit collusion, which in turn substantiates in the way the agents trade at different speeds. The phenomenon is evident thanks to the very low level of asset’s volatility. Thus, evidences of tacit collusive behaviour in this first case exist, moreover the collusive behaviour is adopted by the agents even if they have neither information about the existence nor about the strategy of the other agent trading in the market.

### v2

@@ 3.1.1 Action selection and reward function
For each of the $N$ time-steps and $C$ trading iterations, the agents’ knowledge of the state of market environment is the tuple $g^{i}_{t}=(t,q^{i}_{t},S_{t-1})$ for $i=1,2$. Thus, each agent knows the current time-step, her individual remaining inventory, and the permanently impacted mid-price at the previous time step. In our setting $t\in\{1,\ldots,N\}$ is the discrete time index, $q_{t}\in[0,q_{0}]$ is the agent’s remaining inventory at time $t$, $S_{t-1}\in\mathbb{R}_{+}$ is the (permanently impacted) mid-price at the previous time step. All three components are normalised to lie in [1,1][-1,1] before being used as input to the neural network.
Clearly, only the first and the last are common knowledge of the two agents, while no information on inventory $q_{t}$ or past selling actions $v_{t}$ is shared between the agents.
@@
Then, depending on the current value of $\epsilon$, a draw $\zeta$, different for each agent, from a uniform distribution determines whether the agent performs exploration or exploitation. This means that with probability $\epsilon$ the agent chooses to explore and thus she chooses at time-step $t$ the quantity to sell $v_{t}$ sampling from a normal distribution with mean $\mu=\frac{q_{t}}{N-t}$ and standard deviation $\delta=|\frac{q_{t}}{N-t}|$. Alternatively, with probability $1-\epsilon$, the agent chooses the optimal Q-action as the one that maximises the Q-values from $Q_{\text{main}}$ net, thus exploiting what learnt in the exploration phase. We bind the agents to sell all their inventory within the considered time window, still exploring a large number of states, rewards, and actions.
@@
Once every $m$ actions taken during training iterations by both the agents, $\epsilon$ is multiplied by a constant $c<1$ such that $\epsilon\to 0$ as $m\to\infty$. In this way, for a large number of iterations $C$, $\epsilon$ converges to zero and the algorithm gradually stops exploring and starts to greedily exploit what the agent has learned in terms of weights $\theta$ of the $Q_{\text{main}}$ net. Notice that the update for $\epsilon$ happens at the same rate for both the agents, thus they explore and exploit contemporaneously, while the draw from the uniform distribution $\zeta$ is different for every agent. For each time step $t$ we decide which of the two agents trades first with a coin toss. Once that the ordering of the trades is decided, the action decision rule in the training phase unfolds, for each agent $i=1,2$, as:
@@
$\begin{split}&\epsilon\in(0,1)\,\,,\,\,\zeta_{{i}}\sim\mathcal{U}(0,1)\\ &v^{(i)}_{{t}}=\begin{cases}\sim\mathcal{N}(\mu=\frac{q_{t}}{N-t},\delta=|\frac{q_{t}}{N-t}|)\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,,\text{if}\,\,\zeta\leq\epsilon\\ \operatorname*{arg\,max}_{v^{\prime}\in[{0},q_{0}]}Q_{M}(g^{i}_{t},v^{\prime}|\theta_{\text{main}})\,\,\,\,\,\,\,\,,\text{else}\end{cases}\end{split}$ (22)
The actions chosen following this rule lie on the strategy space for each agent, which is the set of functions:
$\pi:(t,q_{t},S_{t-1})\mapsto v_{t},$
mapping the agent’s state to a trading action. These policies, denoted by $\pi$, are implicitly encoded by the learnt Q-function, since the agent chooses at each state $g_{t}$ the action $v_{t}=\arg\max_{v}Q(g_{t},v;\theta)$. Thus, the strategy space corresponds to the function class that the Q-network can represent and optimise over. Notice that we do not account for the strategy space directly; indeed the Double Deep Q learning algorithm that we developed and used is an off-policy algorithm, thus we do not parametrise the policies but we aim at finding those actions that maximise the Q value which is the output of the Deep Neural Nets (the Q networks), one for each agent. Notice also that the action space for each agent $i$ at each time step $t$ consists of trading volumes set to be $v^{(i)}_{t}\in[0,q^{(i)}_{0}$] Actions are also normalised during training.
@@
[…]
@@
Thus, for each time step $t$ the agent $i$ sees the reward from selling $v_{t,i}$ shares, and stores the state of the environment, $g_{t,i}$, where the sell action was chosen along with the reward and the next state $g_{t+1,i}$ where the environment evolves. At the end of the episode, the reward per episode per agent is $q_{0}S_{0}+\sum^{N}_{t=1}S_{t-1}v_{t,i}-\alpha v_{t,i}^{2}$. Written in this form, the aim of the agent is to cumulatively maximise such reward, so that the liquidation value of the inventory occurs as close as possible to the initial value of the portfolio.
@@ 3.1.2 Training scheme
NN layers 55 $C$ train its. $5,000$ NN intervals $10$
Hidden nodes $30$ $M$ test its. $2,500$ $S_{0}$ price $10\mathdollar$
ADAM lr $0.0001$ mm reset rate $75$ acts. $q_{0}$ inventory $100$
@@
$L$ mem. len. $15,000$ $\gamma$ discount 11 $\kappa$ p. impact $0.001$
Bearing in mind that the procedure is exactly the same for both agents, we focus now on the training scheme of one of them, dropping the subscript $i$. We let the states $g_{t}$, actions $v_{t}$, rewards $r_{t}$ and subsequent future states $g_{t+1}$ obtained by selling a quantity $v_{t}$ in state $g_{t}$, to form a transition tuple that is stored into a memory of maximum length $L$, we have two different memories, one for each agent. As soon as the memory contains at least $b$ transitions, the algorithm starts to train the Q-nets. To this end, the algorithm samples random batches of length $b$ from the memory of the individual agent, and for each sampled transition $j$ they individually calculate:
$y^{j}_{t}(\theta_{\text{tgt}})=\begin{cases}r^{j}_{t}&\text{if }{t}=N;\\ r^{j}_{t}+\gamma Q_{\text{tgt}}(g^{j}_{t+1},v^{*}|\theta_{\text{tgt}})&\text{else}\quad.\end{cases}$ (25)
In Eq. (25), $r^{j}_{t}$ is the reward for the time step considered in the transition $j$, $g^{j}_{t+1}$ is the subsequent state reached in $t+1$, known since it is stored in the same transition $j$, while $v^{*}=\operatorname*{arg\,max}_{v}Q_{\text{main}}(g^{j}_{t},v|\theta_{\text{main}})$. In this context $\gamma$ is a discount factor. Each agent individually minimises the mean squared error loss between the target $y(\theta_{\text{tgt}})$ and the values obtained via the $Q_{\text{main}}$ net. In formulae:
@@
The features of the Q-nets for each agent are ($q_{i,t},t,S_{t-1},v_{i,t}$) and are normalised in the domain [1,1][-1,1] using the procedure suggested in [undefai], whereas normalised mid-prices $\bar{S}_{t}$ are obtained via min-max normalisation.
When considering the inputs to the Q-networks, we choose this set of features for two main reasons. The first one is more theoretical. We want to model the interaction between the agents in such a way that there is no direct exchange of information about the strategy followed by either of them. With this choice of features, the only available ‘shared’ information is the one on the the mid-price level of the traded security, thus no direct communication between the two agents happens. The second reason is more practical. On the one hand we want to be parsimonious with the set of features seen by the agent, and since the environment where the agents trade is a multi-agent Almgren and Chriss model, the inventory and the time step along with the traded quantity, would suffice as features since the problem would be solved with strategies that do not depend on the level of mid-price. However, as shown in [undefai, undefaf] adding additional features that theoretically should be irrelevant, practically help the algorithm learning the optimal trading strategy faster.
In our setup we use fully connected feed-forward neural networks with $5$ layers, each with $30$ hidden nodes. The activation functions are leakyReLu, and finally we use ADAM for optimisation. With the exception of the volatility parameter $\sigma$ which will be specified later, the parameters used in the algorithm are reported in Table 1, the training algorithm is reported in Algorithm 1.
@@
      $v_{i,t}\leftarrow\begin{cases}\text{sample}\,\,\,\mathcal{N}(\mu=\frac{q_{t}}{N-t},\delta=|\frac{q_{t}}{N-t}|)\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,,\text{with probability $\epsilon$}\\ \operatorname*{arg\,max}_{v^{\prime}\in[0,q_{0}]}Q_{M}(g_{t},v^{\prime}|\theta_{\text{main}})\,\,\,\,\,\,\,\,,\text{with probability ($1-\epsilon$)}\end{cases};$
@@
Figure 1: Scatter plot of the IS of the two agents for 20 testing runs of $2,500$ iterations in the zero noise case ($\sigma=10^{-9}$).
@@ 4 Results
In this Section we present the results obtained by using the RL algorithm shown in Algorithm 1 in the game setting outlined in Section 2. The experiments aim at studying the existence and the form of the equilibrium strategies and of their costs, learnt by the agents during the experiments, this is done by analysing the policies chosen by the two interacting trading agents. We then compare the equilibria under the learnt dynamics with the Nash equilibrium strategies, the Pareto-efficient set of solutions, and the Pareto optimal strategy. We consider different scenarios, using the market structure as outlined in Eq. (4).
It can be easily noticed how the theoretical Nash equilibrium for this game does not depend on the volatility level of the asset. In fact, when the risk aversion parameter $\lambda=0$, the Nash equilibrium of Eq. (7) depends only on the permanent and temporary impact coefficients $\kappa$ and $\alpha$, respectively. However, in the numerical determination of the equilibrium solution, volatility $\sigma$ and the associated diffusion play the role of a disturbance term, since an agent cannot determine whether an observed price change is only due to the impact generated by the trading of the other agent or if it is a random fluctuation due to volatility. In this sense, volatility plays the role of a noise-to-signal ratio here. One can interpret the level of $\sigma$ as a measure of presence of noise traders that are randomly affecting the price while the two agents do optimal execution. In this sense, we could also consider $\sigma$ as a measure of market liquidity. To quantify the effect of volatility on the learnt equilibria, we perform three sets of experiments for different levels of the volatility parameter $\sigma$, leaving unchanged the impact parameters. More specifically we consider, for both training and testing phase the case where $\sigma=10^{-9}$ (termed zero noise case), $\sigma=10^{-3}$ (moderate noise) and $\sigma=10^{-2}$ (large noise) case. In all three cases, we use the same temporary and permanent impacts $\alpha=0.002$ and $\kappa=0.001$.
Other preliminary analyses on the behaviour of the agents when one agent follows a Nash or a Pareto strategy, and the other learns a reaction strategy can be found in Appendix A. In what follows we will only concentrate on the simultaneous multi-agent learning problem.
@@ 4.1 The zero noise case
The results of the experiments are displayed in Figure 1, which shows, as a scatter plot, the $IS$ of both agents in each iteration of the testing phase. Each colour represents the results of one of the $20$ runs of $2,500$ iterations each. The blue circles show the $IS$ centroids per testing run (the number in the circle identifies the run). For comparison the plot reports as a red star the point corresponding to the Nash equilibrium of Eq. (7), which is used to divide the graph into four quadrants, delimited by the red dashed lines. The top-right quadrant contains point which are sub-optimal with respect to the Nash equilibrium for both players, while the points in the top-left and bottom-right quadrant report points where the found equilibrium favours one of the two agents at the expenses of the other. In particular, in these regions one of the agents achieve an IS smaller than the one of the Nash equilibrium, while the other agents performs worst. In a sense, one of the agents predates on the other in terms of reward. The bottom-left quadrant is the most interesting, since here both agents are able to achieve an $IS$ smaller than the one in the Nash equilibrium and thus contain supra-competitive $IS$ costs that are potential collusive equilibria. For comparison, the black star indicates the $IS$ of the Pareto-optimal strategy found in Theorem 2, while the magenta line is the linear combination of the Pareto-optimal $IS$. Finally, the green rectangle denotes the area between the Nash and the Pareto-optimal $IS$s, and we call it the supra-competitive area. In fact, in that area we find the $IS$s for both the agents that lie between the collusion defined by the Pareto-optimum and the Nash equilibrium. It is the area where, although the $IS$ costs achieved do not overlap perfectly with the Pareto optimal one, thus the collusion, are still supra-competitive.
@@
Figure 2: Optimal execution strategies for 20 testing runs of $2,500$ iterations, using $\sigma=10^{-9}$.
@@
Looking at Figure 1 we can notice that the $IS$s per testing run concentrate in the supra-competitive area of the graph (green rectangle) between the Nash and the Pareto-optimal $IS$. This means that when the noise is minimal, it is easier for the agents to adopt strategies whose cost levels are supra-competitive and potentially consistent with tacit collusive strategies. The agents are able to obtain costs that fall near the collusion $IS$, i.e. the Pareto-optimum. The majority of the remaining $IS$s per iteration still lie within the second and fourth quadrant of the graph, meaning that again they are in general able to find strategies that, at each iteration, do not allow one agent to be better without worsening the other.
@@
Considering the strategies found by the agents in Figure 2, we notice how the agents trade at different speeds, i.e. their selling policies are consistent with the presence of a slow trader and a fast trader. It can be seen how the policies followed by the agents in almost all the 20 simulations depend inversely on the strategy adopted by the competitor. Moreover, in most of the runs in the supra-competitive region, the average strategy of the agents is very similar to the TWAP strategy. This is possible thanks to very low noise, and in this case the agents are able to find more easily the Pareto-optimal strategy, which corresponds to an equilibrium where they both pay the lowest amount possible in terms of $IS$.
@@
These evidences underline the quite intuitive fact that, when the noise-to-signal ratio is low it is simple for the agents to disentangle their actions from those of other agents, thus making it easier for the agents to learn and possibly infer the strategy of the competitor. The agents find an incentive to deviate from the Nash-equilibrium adopting strategies that allow for lower costs. The majority of these strategies correspond to an average supra-competitive $IS$, thus with lower cost than in the Nash equilibrium, but still slightly greater than the Pareto-optimum in terms of $IS$. In general, per iteration neither agent can be better off without increasing the other agent cost level.
Coupling the cost plot in Figure 1 with the average selling schedules in Figure 2, it can be seen how the centroids that relate to higher costs for one of the agents correspond to a deviation from the TWAP strategy by either of them, such deviation corresponds to another deviation by the other agent, that substantiates in a faster selling rate. There is, on the one hand, evidence of supra-competitive behaviour, on average, when $IS$ costs are considered, which in turn is reflected in the way the agents trade at different speeds. The phenomenon is evident thanks to the very low level of asset’s volatility. Thus, hinting to tacit collusive behaviour in this first case, judging by the supra-competitive cost outcomes. In Section 5 we thoroughly discuss this point.